y=x^2e^x find the indefinite integral funtion



1
Expert's answer
2021-03-04T10:40:17-0500

Let's integrate "\int x^2e^xdx" by parts. Let's use the rule:


"\int fg'=fg-\int f'g,"

here, "f=x^2, g=e^x".

Then, we get:


"\int x^2e^xdx=x^2e^x-2\int xe^xdx."

Let's integrate "\int xe^xdx" by parts and use the same rule. This time "f=x, g=e^x". Finally, we get:


"\int x^2e^xdx=x^2e^x-2(xe^x-\int e^xdx),""\int x^2e^xdx=x^2e^x-2xe^x+2e^x=(x^2-2x+2)e^x+C."